# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ 3.1. Matrix representation of $F_{w}$, boundedness and a perturbation result
This result follows from the above matrix representation of $F_{w}$. Consider the similarity operator $V:\ell^{p}_{a,b}(\mathtt{A})\rightarrow\ell^{p}(\mathbb{Z})$ given by $V\big{(}\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\big{)}=\sum_{n=-\infty}^{\infty}\lambda_{n}e_{n}$, where $\{e_{n}\}_{n\in\mathbb{Z}}$ is the standard basis in $\ell^{p}(\mathbb{Z})$. Now, from the proof of Theorem 3.4 we understand that $F_{w}$ acting on $\ell^{p}_{a,b}\mathtt{(}A)$ is similar via $V$ to the operator given by the series (which is convergent absolutely with respect to the operator norm)
@@
Further, by our assumptions, the operators $E_{-i},$ $D$ and $E_{i}$ are compact on $\ell^{p}(\mathbb{Z})$ as the entries in each of the matrices $E_{-i},$ $D$ and $E_{i}$ converge to 00 except for the case $i=1$. The later case corresponds to $E_{1}$ which is the weighted forward shift on $\ell^{p}(\mathbb{Z})$ with weights $\alpha=\{\alpha_{n}\}_{n=-\infty}^{\infty}$, as in the theorem. We take $E_{\alpha}:=E_{1}$, and $K:=\sum_{i=1}^{\infty}E_{-i}+D+\sum_{i=2}^{\infty}E_{i}$ which is a compact operator on $\ell^{p}(\mathbb{Z})$. This proves the theorem. ∎
@@ 3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{1}{w_{\nu}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert w_{\nu-1}\cdots w_{\nu-n}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert,$ (3.2)
@@
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{1}{w_{\nu}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert w_{\nu-1}\cdots w_{\nu-n}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert,$ (3.3)
@@
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert=\infty,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.4)
@@
$s(z^{\nu})=\frac{1}{w_{n-1}}z^{\nu-1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
@@
A necessary condition for the supercyclicity of $F$ is stated below.
Proposition 3.8.
Let $X$ be a Banach space of analytic functions over an annulus $r<|z|<R$, having bounded evaluation functionals. Let the unweighted forward shift $F$ be bounded on $X$. Then $F$ cannot be supercyclic.
Proof.
Since $F$ is bounded on $X$, we have
$f\in X\Rightarrow zf\in X.$
Thus, $F$ is the multiplication operator $M_{\varphi}(f)=\varphi f$ defined on $X$, where $f\in X$, and $\varphi(z)=z$. However, multiplication operators cannot be supercyclic as its adjoint has more than two distinct eigenvalues with eigenvectors being the non-zero evaluation functionals, cf. [20]. ∎
@@
    $\sum_{n=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{n-1}}{a_{n}}\right\rvert^{p}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\left\lvert\frac{1}{w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}}\right\rvert^{p}<\infty.$
@@
    For $\nu\in\mathbb{N}$,
    $\sum_{n=1}^{\infty}\left\lvert w_{\nu}\cdots w_{\nu+n-1}\frac{a_{0}\cdots a_{\nu+n-1}}{b_{0}\cdots b_{\nu+n-1}}\right\rvert^{p}<\infty~{}~{}\text{and}~{}~{}\displaystyle\sum_{n=1}^{\infty}\left\lvert\frac{1}{{w_{\nu-1}\cdots w_{\nu-n}}}\frac{b_{-1}\cdots b_{\nu-n}}{a_{-1}\cdots a_{\nu-n}}\right\rvert^{p}<\infty.$
Then $(i)\Rightarrow(ii)\Rightarrow(iii).$ Also, the conditions $(iii)$ and $(iv)$ together yield $(i)$.
@@
$\cdots+w_{-km-1}w_{-km}\cdots w_{-2}A_{-km-1}\frac{1}{z}+w_{-km}w_{-km+1}\cdots w_{-1}A_{-km}+\cdots$ $=$
$\ldots+A_{2}z^{-2}+A_{-1}z^{-1}+A_{0}+\cdots+A_{n}z^{n}\cdots,$
@@
for all $z\in\mathtt{A}$. Using the coefficients of like powers for comparison, we obtain $A_{j+km}=w_{j}w_{j+1}\cdots w_{j+km-1}A_{j}$, $\forall~{}~{}k\in\mathbb{N}$, where $10\leq j\leq m-$ 1. For $j=0,$ we have $\lambda_{km}a_{km}+\lambda_{km-1}b_{km-1}=w_{0}w_{1}\cdots w_{km-1}(\lambda_{0}a_{0}+\lambda_{-1}b_{-1})~{}\forall~{}~{}k\geq 1.$ Hence, we get
$\lvert\lambda_{0}a_{0}+\lambda_{-1}b_{-1}\rvert^{p}\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{km-1}}{a_{km}}\right\rvert^{p}\leq C_{1}\left(r^{p}\sum_{k=1}^{\infty}\left\lvert\lambda_{km-1}\right\rvert^{p}+\sum_{k=1}^{\infty}\left\lvert\lambda_{km}\right\rvert^{p}\right),$
for some $C_{1}>0$. Here, $r:=\sup_{n\geq 0}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert$. As $\{\lambda_{n}\}\in\ell^{p}(\mathbb{Z})$, we have
$\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{km-1}}{a_{km}}\right\rvert^{p}<\infty.$
Along the same lines, for $j=1,2,\ldots,m-1,$ we have $\lambda_{j+km}a_{j+km}+\lambda_{j+km-1}b_{j+km-1}=\\ w_{j+1}w_{j+2}\cdots w_{j+km}(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1})$, and
$\sum_{k=1}^{\infty}\left\lvert\frac{w_{j}w_{j+1}\cdots w_{j+km-1}}{a_{j+km}}\right\rvert^{p}<\infty.$
Consequently, the series $\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{k-1}}{a_{k}}\right\rvert^{p}$ is convergent. In a similar way, the series
@@
$\sum_{k=1}^{\infty}\left\lvert\frac{1}{w_{-1}w_{-2}\cdots w_{-k}b_{-k-1}}\right\rvert^{p}$ can be shown to be convergent, and thus the implication $(ii)\Rightarrow(iii)$ follows.
We now prove that $(iv)$ implies $(i)$ by applying the chaoticity criterion. Take $D$ to be the space of all Laurent polynomials, and define $S:D\rightarrow D$ by $S(z^{\nu})=\frac{1}{w_{\nu-1}}z^{\nu-1},~{}~{}\nu\in\mathbb{Z}.$ Clearly, $F_{w}S=I$ on $D.$ It remains to show that the series $\sum_{n=0}^{\infty}F_{w}^{n}(f)$ and $\sum_{n=0}^{\infty}S^{n}(f)$ converge unconditionally. It suffices to prove that
$\sum_{n=1}^{\infty}w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}$
are unconditionally convergent in $\ell^{p}_{a,b}(\mathtt{A}),$ for all $\nu\in\mathbb{N}$. Recalling the Schauder basis expansion from (2.3) and using the Proposition 2.4, for a fixed $\nu\in\mathbb{N}$, we obtain
$\left\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\right\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\leq M_{1}\left\lvert\frac{(w_{\nu}\cdots w_{\nu+n-1})(a_{0}\cdots a_{\nu+n-1})}{b_{0}\cdots b_{\nu+n-1}}\right\rvert$
and
$\left\lVert\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}\right\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\leq M_{2}\left\lvert\frac{b_{-1}\cdots b_{\nu-n}}{(w_{\nu-1}\cdots w_{\nu-n})(a_{-1}\cdots a_{\nu-n})}\right\rvert,$
for some constants $M_{1}$ and $M_{2}$. It follows that the right hand side of the above inequality is $p$-summable. Consequently, the series $\sum_{n=1}^{\infty}w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}$ are convergent. The unconditional convergence follows because $\{f_{n}\}$ is a normalized Schauder basis for $\ell^{p}_{a,b}(\mathtt{A})$ which is invariant under $\ell^{1}$ convolutions. We can now conclude that $F_{w}$ satisfies the chaoticity criterion, and (i) follows. ∎
@@
    $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert=0=\liminf_{n\rightarrow\infty}~{}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.5)
@@
    $F_{w}^{*}$ is topologically mixing on $\ell^{p}_{a,b}(\mathtt{A})^{*}$ if
    $\lim_{n\rightarrow\infty}\lvert~{}w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert=0=\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.6)
@@
    $\liminf_{n\rightarrow\infty}\lvert~{}w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.7)
@@
Suppose that (3.5) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}^{*}$ is hypercyclic by taking $D=\text{span}\{k_{n}:n\in\mathbb{Z}\}$, and $S$ as the map defined by
$S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$
By Proposition 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that $(F_{w}^{*})^{n_{k}}(k_{\nu})$ and $S^{n_{k}}(k_{\nu})$ both tends to 0 as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
@@
for some constants $M_{1},M_{2}>0$ and $n_{k}>\nu.$ Now by the assumption (3.5), we get that $F_{w}^{*}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})^{*}.$
@@
Let $\ell^{p}_{a,b}(\mathtt{A})$ $1<p<\infty,$ be an analytic Banach space over the unit annulus corresponding to $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfying the conditions of Theorem 3.4. Consider the following statements:

### v2

@@ 3.1. The matrix of $F_{w}$, boundedness, compact perturbations, and the essential spectrum
This result follows from the above matrix representation of $F_{w}$. Consider the similarity operator $V:\ell^{p}_{a,b}(\mathtt{A})\rightarrow\ell^{p}(\mathbb{Z})$ given by $V\big{(}\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\big{)}=\sum_{n=-\infty}^{\infty}\lambda_{n}e_{n}$, where $\{e_{n}\}_{n\in\mathbb{Z}}$ is the standard basis in $\ell^{p}(\mathbb{Z})$. Now, from the proof of Theorem 3.3 we understand that $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is similar via $V$ to the operator given by the series (which is convergent absolutely with respect to the operator norm)
@@
Further, by our assumptions, the operators $E_{-i},$ $D$ and $E_{i}$ are compact on $\ell^{p}(\mathbb{Z})$ as the entries in each of $E_{-i},$ $D$ and $E_{i}$ converge to 00 except for the case $i=1$. Also, the operator $E_{1}$ is the weighted forward shift on $\ell^{p}(\mathbb{Z})$ with weights $\alpha=\{\alpha_{n}\}_{n=-\infty}^{\infty}$, as in the statement of this theorem. We now take
$E_{\alpha}:=E_{1}$, and $K:=\sum_{i=1}^{\infty}E_{-i}+D+\sum_{i=2}^{\infty}E_{i}$,
required as in the theorem since $K$ is a compact operator on $\ell^{p}(\mathbb{Z})$.
Since the essential spectrum is invariant under compact perturbations, we have $\sigma_{e}(F_{w})=\sigma_{e}(E_{\alpha}+K)=\sigma_{e}(E_{\alpha})$. ∎
@@ 3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
    $\limsup_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{\nu+k}^{-1}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{\nu-k}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert,$ (3.2)
@@
    $\lim_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{\nu+k}^{-1}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}\right\rvert=\infty=\lim_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{\nu-k}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert$ (3.3)
@@
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{\prod_{k=1}^{n}w_{\nu-k}}{\prod_{k=0}^{n-1}w_{\nu+k}}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}~{}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert=\infty,\hskip 14.22636pt\forall~{}~{}\nu\in\mathbb{N}.$ (3.4)
@@
$S(z^{\nu})=\frac{1}{w_{\nu-1}}z^{\nu-1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
@@
Remark 3.8.
It should be noted that the unweighted forward shift $F$ cannot be supercyclic on a Banach space $\mathcal{E}$ of analytic functions over an annulus $r<|z|<R$, having bounded evaluation functionals. Indeed, the boundedness of $F$ on $\mathcal{E}$ is equivalent to the following: $f\in\mathcal{E}\Rightarrow zf\in\mathcal{E}.$ Thus, $F$ is the multiplication operator $M_{\varphi}(f)=\varphi f$ defined on $\mathcal{E}$, where $f\in\mathcal{E}$, and $\varphi(z)=z$. However, a multiplication operator cannot be supercyclic as its adjoint has more than two distinct eigenvalues with eigenvectors being the non-zero evaluation functionals, cf. [19].
@@
    $\sum_{n=1}^{\infty}\left\lvert\frac{1}{a_{n}}\prod_{k=0}^{n-1}w_{k}\right\rvert^{p}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\left\lvert\frac{1}{b_{-n-1}}\frac{1}{\prod_{k=1}^{n}w_{-k}}\right\rvert^{p}<\infty.$
@@
    $\sum_{n=1}^{\infty}\left\lvert\prod_{k=0}^{n-1}\frac{a_{k}w_{k}}{b_{k}}\right\rvert^{p}<\infty~{}~{}\text{and}~{}~{}\displaystyle\sum_{n=1}^{\infty}\left\lvert\prod_{k=1}^{n}\frac{b_{-k}}{w_{-k}a_{-k}}\right\rvert^{p}<\infty.$
Then $(i)\Rightarrow(ii)\Rightarrow(iii)$ and $(iv)\Rightarrow(iii)$ Also, the condition $(iv)\Rightarrow(i).$
@@
$\sum_{n=-\infty}^{\infty}w_{n}w_{n+1}\cdots w_{n+km-1}A_{n}z^{n+km}=\sum_{n=-\infty}^{\infty}A_{n}z^{n},$
@@
for all $z\in\mathtt{A}$. Using the coefficients of like powers for comparison, we obtain $A_{j+km}=w_{j}w_{j+1}\cdots w_{j+km-1}A_{j}$, $\forall~{}~{}k\geq 1$, where $10\leq j\leq m-$ 1. For $j=0,$ we have $\lambda_{km}a_{km}+\lambda_{km-1}b_{km-1}=w_{0}w_{1}\cdots w_{km-1}(\lambda_{0}a_{0}+\lambda_{-1}b_{-1})$, for $k\geq 1.$ Writing $s_{n,j}:=w_{j}w_{j+1}\cdots w_{j+n-1}$ for convenience, we see that
$\lvert\lambda_{0}a_{0}+\lambda_{-1}b_{-1}\rvert^{p}\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,0}}{a_{km}}\right\rvert^{p}\leq C_{1}r^{p}\big{(}\sum_{k=1}^{\infty}\left\lvert\lambda_{km-1}\right\rvert^{p}+\sum_{k=1}^{\infty}\left\lvert\lambda_{km}\right\rvert^{p}\big{)},$
@@
for some constant $C_{1}>0$. Here, $r:=\sup_{n\geq 0}\left\lvert b_{n}/a_{n+1}\right\rvert$. As $\{\lambda_{n}\}\in\ell^{p}(\mathbb{Z})$, we have $\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,0}}{a_{km}}\right\rvert^{p}<\infty.$ Along the same lines, for $j=1,2,\ldots,m-1,$ we have $\lambda_{j+km}a_{j+km}+\lambda_{j+km-1}b_{j+km-1}=w_{j}w_{j+1}\cdots w_{j+km-1}(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1})$, and $\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,j}}{a_{j+km}}\right\rvert^{p}<\infty.$ Consequently, we get the convergence of $\sum_{k=1}^{\infty}\left\lvert\frac{s_{k,0}}{a_{k}}\right\rvert^{p}$. In a similar way, it can be proved that $\sum_{k=1}^{\infty}\left\lvert w_{-1}w_{-2}\cdots w_{-k}b_{-k-1}\right\rvert^{-p}$ is convergent, and thus the implication $(ii)\Rightarrow(iii)$ follows.
The implications in $(iv)\Rightarrow(iii),$ are easily seen to hold, just by noting the expressions
$\frac{1}{a_{n}}=\frac{a_{0}\cdots a_{n-1}}{b_{0}\cdots b_{n-1}}\frac{b_{0}\cdots b_{n-1}}{a_{0}\cdots a_{n}}~{}\text{and}~{}\frac{1}{b_{-n-1}}=\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n}}\frac{a_{-1}\cdots a_{-n}}{b_{-1}\cdots b_{-n-1}},$
along with $\limsup_{n}|b_{n}/a_{n+1}|<1$, and $\limsup_{n}|a_{-n+1}/b_{-n}|<1$.
We now prove that $(iv)$ implies $(i)$ by applying the chaoticity criterion. Take $D$ and $S$ to be the same set and map as in the previous theorem. It suffices if we show that
$\sum_{n=1}^{\infty}w_{0}w_{1}\cdots w_{n-1}z^{n}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}z^{-n}$
are unconditionally convergent in $\ell^{p}_{a,b}(\mathtt{A}).$ Recalling the expansion of $z^{n}$ in terms of the basis $\{f_{j}\}$, and to keep calculations simple, let us assume that $s_{n}=w_{0}w_{1}\cdots w_{n}$ and $u_{n}=\frac{a_{0}\cdots a_{n}}{b_{0}\cdots b_{n}}$ for all $n\geq 0$ and we obtain
$\sum_{n=1}^{\infty}s_{n-1}z^{n}$ $=$ $\frac{1-\lambda_{0}a_{0}}{b_{-1}}\left(\sum_{j=0}^{\infty}(-1)^{j+1}s_{j}u_{j}\right)\times\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+\frac{\lambda_{0}a_{0}-1}{a_{0}}\left(\sum_{j=0}^{\infty}(-1)^{j+1}s_{j}u_{j}\right)f_{0}+\frac{\lambda_{0}b_{0}}{a_{1}}\left(s_{0}u_{0}+\frac{\lambda_{0}a_{0}-1}{a_{0}}\sum_{j=1}^{\infty}(-1)^{j}s_{j}u_{j}\right)f_{1}$
$+\sum_{j=2}^{\infty}\left(\frac{\lambda_{0}a_{0}}{b_{j}u_{j}}\left(\sum_{k=1}^{j}(-1)^{k-1}s_{j-k}u_{j-k}+\frac{\lambda_{0}a_{0}-1}{a_{0}}\sum_{k=0}^{\infty}(-1)^{k+1}s_{j+k}u_{j+k}\right)\right)f_{j}.$
Now, using the conditions in (iv) we can show that the coefficients in the previous expansion in terms of $\{f_{n}\}$ form a member in $\ell^{p}(\mathbb{Z})$, and hence, it converges unconditionally as $\{f_{n}\}$ is an unconditional basis. Thus, series $\sum_{n=1}^{\infty}w_{0}\cdots w_{n-1}z^{n}$ is convergent. It remains to show that $\sum_{n=1}^{\infty}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}z^{-n}$ is unconditionally convergent, for that we will write the expansion of $z^{-n}$ in terms of the basis $\{f_{j}\}$ and then proceed as above. We can now conclude that $F_{w}$ satisfies the chaoticity criterion, and (i) follows. ∎
@@
    $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\left\lvert b_{\nu-n-1}\prod_{k=1}^{n}w_{\nu-k}\right\rvert=0=\liminf_{n\rightarrow\infty}~{}\left\lvert\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.5)
@@
    $F_{w}^{*}$ is topologically mixing on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\lim_{n\rightarrow\infty}\left\lvert b_{\nu-n-1}\prod_{k=1}^{n}w_{\nu-k}\right\rvert=0=\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.6)
@@
    $\liminf_{n\rightarrow\infty}\left|b_{\nu-n-1}a_{\nu+n}\frac{\prod_{k=1}^{n}w_{\nu-k}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right|=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.7)
@@
Suppose that (3.5) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}^{*}$ is hypercyclic by taking $D=\text{span}\{k_{\nu}:\nu\in\mathbb{Z}\}$, and $S$ as the map defined by
$S(k_{\nu})=\frac{1}{w_{\nu}}k_{\nu+1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
The map $S$ is well-defined as $\{k_{\nu}:\nu\in\mathbb{Z}\}$ is linearly independent. Indeed, if $\sum_{\nu=m_{1}}^{m_{2}}\alpha_{\nu}k_{\nu}=0$ for some $m_{1}<m_{2}$, then $\alpha_{\nu}=0$ because of the following: $k_{\nu}(z^{m})=0$ if and only if $\nu\neq m$ and $k_{\nu}(z^{\nu})\neq 0$.
By Proposition 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that, for some increasing $(n_{k})\subset\mathbb{N},$ the sequences $(F_{w}^{*})^{n_{k}}(k_{\nu})$ and $S^{n_{k}}(k_{\nu})$ both tend to 0 as $k\to\infty,$ $\nu\in\mathbb{N}.$ But this is clear by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
@@
for some constants $M_{1},M_{2}>0$. Now by the assumption (3.5), we get that $F_{w}^{*}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$
@@
Our next aim is to obtain necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic. We need a representation of the functional $f_{n}^{*}$ in terms of $k_{n}$.
Proposition 3.11.
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\widehat{f}(n),$ defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. The following hold.
  • (i)
    For $n\geq 1,$
    $f_{n}^{*}=\frac{1}{a_{n}}k_{n}+\sum_{j=1}^{n}(-1)^{j}\frac{b_{n-1}b_{n-2}\cdots b_{n-j}}{a_{n}a_{n-1}\cdots a_{n-j}}k_{n-j}+(-1)^{n+1}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}f_{-1}^{*}.$
  • (ii)
    For $n\geq 2,$
    $f_{-n}^{*}=\frac{1}{b_{-n}}k_{-n+1}+\sum_{j=2}^{n}(-1)^{j-1}\frac{a_{-n+1}a_{-n+2}\cdots a_{-n+j-1}}{b_{-n}b_{-n+1}\cdots b_{-n+j-1}}k_{-n+j}+(-1)^{n}\frac{a_{-n+1}\cdots a_{0}}{b_{-n}\cdots b_{-1}}f_{0}^{*}.$
Proof.
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. Rearranging into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ Now for $f\in\ell^{p}_{a,b}(\mathtt{A}),$ we have
$\frac{1}{a_{n}}k_{n}(f)-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0}(f)$ $=$ $\lambda_{n}+(-1)^{n}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}\lambda_{-1}$
$=$ $f_{n}^{*}(f)+(-1)^{n}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}f_{-1}^{*}(f),$
for all $n\geq 1$, which gives (i).
Proof of (ii) is similar. ∎
@@
Let $\ell^{p}_{a,b}(\mathtt{A}),$ $1<p<\infty,$ be an analytic Banach space over the unit annulus corresponding to $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfying the conditions of Theorem 3.3. Consider the following statements: